An ionization chamber with parallel conducting plates as anode and cathode has 5 × 107 electrons and the same number of singly charged positive ions per cm3. The electrons are moving towards the anode with velocity 0.4 m/s. The current density from anode to cathode is 4 mA/m2. The velocity of positive ions moving towards cathode is


        (a)    0.4 m/s     


(b)    1.6 m/s


        (c)    Zero


        (d)    0.1 m/s


Grade:11

1 Answers

Amit Saxena
10 years ago

(d)

Current = Ie + Ip.

Ie and Ip are current due to electrons and positively charged ions.

I = ne AVd

Given, 1/A = 4 × 10-6 A/m2

4 × 10-6 ×A = 5 × 10-6 × 1.6 × A(v+0.4)

 = v + 0.4 Þ 0.5 = v + 0.4 Þ v = 0.1 m/s